OpenStudy (anonymous):

Write the exponential expression 3x^3/8 in radical form.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Ex: X^1/2 ___________ The numerator (1) is the power. The denominator (2) is the root. Radical form: \[\sqrt{X}\]

OpenStudy (anonymous):

Ex : X^3/5 The numerator (3) is the power. The denominator (5) is the root. Radical form :\[\sqrt[5]{X^3}\]

OpenStudy (anonymous):

so for mine it would be 8 root 3x^3?

OpenStudy (anonymous):

Yours : 3x^3/8 The numerator (3) is the power. The denominator (8) is the root Radical form : what do u think ?

OpenStudy (anonymous):

its \[\sqrt[8]{3x^3}\]

OpenStudy (anonymous):

Your are right ^_^

OpenStudy (anonymous):

yea thats what i got

OpenStudy (anonymous):

Good job xD